OpenStudy (anonymous):

help please what the equation of tangent line a curve x^3+2x^2+1 absis=1

OpenStudy (calculusfunctions):

Find the equation of the tangent line to the curve\[y =x ^{3}+2x ^{2}+1\]Is this your question?

OpenStudy (calculusfunctions):

at x = 1?

OpenStudy (calculusfunctions):

To find the equation of a tangent to a curve y = f(x) at x = a: i). Find y when x = a. In other words, find f(a) so that you have the point (a, f(a)) ii). Evaluate the derivative (slope of a tangent) at x = a. iii). Use the slope from step ii) and the point (a, f(a)) from step i) to find the equation of the tangent y = mx + b.

OpenStudy (anonymous):

y=x^3+2x^2+1 at x=1 y=1+2+1 y=4 slope dy/dx=3x^2+2x at x=1 dy/dx=3+2=5 slope=5 equat of tangent y-4=5(x-1)